# Terminal 3-folds that are not Cohen-Macaulay

### v1

@@ 3 Review of ramification theory
Let $G=\text{\bf Z}/p=\langle\sigma:\sigma^{p}=1\rangle$ act nontrivially on a normal noetherian integral scheme $Y$. Assume that $Y$ is of finite type over a field or over $\text{\bf Z}_{p}$, so that we can talk about the canonical class $K_{Y}$. Write $f$ for the quotient map $Y\rightarrow Y/G$. For each irreducible divisor $E$ in $Y$ that is mapped into itself by $G$, let $F$ be its image (as an irreducible divisor) in $Y/G$. Assume that $p=0$ on $E$. There are several invariants we want to compute in this situation: the ramification index of the divisor $E$ in $Y$ (the positive integer $e$ such that $f^{*}F=eE$), and the coefficient $c$ of $E$ in the ramification divisor (meaning that $K_{Y}=f^{*}K_{Y/G}+cE$ near the generic point of $E$). Another name for $c$ is the valuation of the different $v_{E}(\mathcal{D}_{L/K})$, where $L$ is the function field $k(Y)$, $K=k(Y/G)$, and the valuation $v_{E}$ is the order of vanishing along $E$. Here $ef=p$, where $f$ is the degree of the field extension $k(E)$ over $k(F)$.
@@
In the ramified case, the relative canonical class (that is, the valuation of the different) can be computed as follows. Define the Artin ramification number $i(E)$ of $Y$ over $Y/G$ along $E$ to be the coefficient of $E$ in the fixed point scheme $Y^{G}$. Equivalently, in terms of $I(a)=\sigma(a)-a$:
@@
Then the valuation of the different is $(p-1)i(E)$ [27, section 2.1]. When $Y$ is regular, so that $Y/G$ is Q-factorial, we can equivalently say that
@@
In the ramified case, the field $k(E)$ is purely inseparable over $k(F)$, with degree $f$ equal to 1 or $p$. We can distinguish the two cases as follows. Since $ef=p$, either $e=p$ and $f=1$ (called wild ramification) or $e=1$ and $f=p$ (called fierce ramification). Let $t$ be a defining function of $E$, that is, a rational function on $Y$ with valuation $v_{E}(t)=1$. It is clear that $v_{E}(I(t))\geq i(E)$. A very convenient criterion is: $Y\to Y/G$ is wildly ramified along $E$ if and only if equality holds [27, section 2.1]. Otherwise, it is fiercely ramified.
@@ 4 Toric divisors
    Let $G=\text{\bf Z}/p$ act on a regular scheme $U$, with the assumptions of Theorem 2.2. Assume moreover that the function $e$ is the greatest common divisor of the functions $I(x_{i})$ for $i=1,\ldots,n$ in the local ring $O_{U,P}$. Then, for each $y\in\operatorname{\mathfrak{m}}_{U}$ such that $\{y=0\}$ is an irreducible divisor $D\subset U$ and $I(y)\in(ey)$, $D$ is the pullback of a Weil divisor $F$ in $U/G$.
@@
    Continue to assume that $I(x_{i})\in e\operatorname{\mathfrak{m}}_{U}$ for $i=1,\ldots,n$, so that $U/G$ is the quotient of a regular scheme $Q$ by $\mu_{p}$. Let $y_{1},\ldots,y_{r}$ be functions on $U$, vanishing at $P$, that are linearly independent in $\operatorname{\mathfrak{m}}_{U}/\operatorname{\mathfrak{m}}_{U}^{2}$. Suppose that $I(y_{j})\in(ey_{j})$ for $j=1,\ldots,r$. Then the corresponding $\mu_{p}$-eigenfunctions $h_{1},\ldots,h_{r}$ on $Q$ (from (2)) are linearly independent in $\operatorname{\mathfrak{m}}_{Q}/\operatorname{\mathfrak{m}}_{Q}^{2}$. That is, these functions are part of a toric coordinate system on $Q$. Finally, if $I(s)=es$ in Theorem 2.2 (as we can assume), then the $\mu_{p}$-weight of $h_{j}$ is equal to the eigenvalue of $\varphi$ on $y_{j}\in\operatorname{\mathfrak{m}}_{U}/\operatorname{\mathfrak{m}}_{U}^{2}$, which is in $\text{\bf F}_{p}$.
@@
(1) Let $E$ be the irreducible divisor $\{y=0\}$ on $U$. The assumptions imply that $E$ is mapped into itself by $G$, and that there is an $i\in\{1,\ldots,n\}$ such that $v_{E}(I(y))>v_{E}(I(x_{i}))$. By section 3, $f\colon U\to U/G$ is either unramified (if $E$ is not fixed by $G$) or fiercely ramified along $E$. In either case, there is an irreducible divisor $F$ on $U/G$ such that $E=f^{*}F$, as we want. (The divisor $F$ need not be Cartier, but the pullback of a Weil divisor is still a Weil divisor (with integer coefficients). Indeed, $F$ is a Cartier divisor outside a codimension-2 subset of $U/G$, by normality of $U/G$.)
@@
(2) Since $Q$ is regular, the pullback of $F$ to $Q$ is a Cartier divisor, hence (after possibly shrinking $U$ and $Q$) of the form $\{t=0\}$ for some function $t\in\operatorname{\mathfrak{m}}_{Q}-\{0\}$. Clearly the divisor $\{t=0\}$ is $\mu_{p}$-invariant. I claim that $t$ times some unit is a $\mu_{p}$-eigenfunction on $Q$. Indeed, in algebraic terms, the action of $\mu_{p}$ on $Q$ makes $O(Q)$ a comodule over $O(\mu_{p})$, and the ideal $(t)$ is an sub-$O(\mu_{p})$-comodule. Every $O(\mu_{p})$-comodule (with no finiteness assumption needed) is the direct sum of its weight spaces, indexed by $\text{\bf Z}/p$. So we can write $t=t_{0}+\cdots+t_{p-1}$ with $t_{i}\in(t)$ and $t_{i}$ of weight $i$. Since $t_{i}\in(t)$, we can write $t_{i}=a_{i}t$ for some $a_{i}\in O_{Q,P}$. Since $O_{Q,P}$ is regular, it is a domain, and hence $1=a_{0}+\cdots+a_{p-1}$. So at least one $a_{i}$ is not in $\operatorname{\mathfrak{m}}_{Q}$, hence is a unit. Then $h:=t_{i}=a_{i}t$ is a unit times $t$ and also a $\mu_{p}$-eigenfunction (of weight $i$), as we want.
@@
(3) After multiplying $e$ by a unit, we can assume that $I(s)=es$, in the terminology of Theorem 2.2. The assumption that $I(y_{j})\in(ey_{j})$ for $j=1,\ldots,r$ implies that $y_{1},\ldots,y_{r}$ in $\operatorname{\mathfrak{m}}_{U}/\operatorname{\mathfrak{m}}_{U}^{2}$ are eigenfunctions of the map $\varphi$. By the proof of Theorem 2.2, the corresponding eigenvalues are in $\text{\bf F}_{p}\subset k_{U}$.
@@
For $j=1,\ldots,r$, we know from (1) that the divisor $\{y_{j}=0\}$ on $U$ is pulled back from a divisor $F$ on $U/G$. (Here $F$ is a Weil divisor, but it is a Cartier divisor outside a codimension-2 subset of $U/G$, since $U/G$ is normal.) By (2), $F$ pulls back to a divisor $\{h_{j}=0\}$ on $Q$ with $h_{j}$ a $\mu_{p}$-eigenfunction. It follows that the Cartier divisors $\{y_{j}=0\}$ on $U$ and $\{h_{j}=0\}$ on $Q$ have the same pullback to $W$; that is, $h_{j}=y_{j}(\text{unit})$ on $W$. Since $h_{j}$ is a $\mu_{p}$-eigenfunction of some weight $b_{j}\in\{0,\ldots,p-1\}$, we have $h_{j}=g_{j}w^{b_{j}}$ for some function $g_{j}$ on $U$, in the notation of the proof of Theorem 2.2. Therefore, $g_{j}=y_{j}(\text{unit})$ on $U$. Since $y_{1},\ldots,y_{r}$ are linearly independent in $\operatorname{\mathfrak{m}}_{U}/\operatorname{\mathfrak{m}}_{U}^{2}$, the same is true for $g_{1},\ldots,g_{r}$. By the proof of Theorem 2.2, the $\mu_{p}$-weight of $h_{j}$ is equal to the eigenvalue of $\varphi$ on the eigenvector $g_{j}\in\operatorname{\mathfrak{m}}_{U}/\operatorname{\mathfrak{m}}_{U}^{2}$. Since $g_{j}\in\operatorname{\mathfrak{m}}_{U}/\operatorname{\mathfrak{m}}_{U}^{2}$ is a nonzero multiple of $y_{j}$, this is the same as the eigenvalue of $\varphi$ on $y_{j}$.
@@ 5 The example over the 2-adic integers
The scheme $Y$ is regular, being an open subset of the affine plane over the discrete valuation ring $\text{\bf Z}_{2}[\mu_{4}]=\text{\bf Z}_{2}[i]/(i^{2}+1)$. Since $Y$ is a normal integral affine scheme of dimension 3, so is $Y/G$. The ring $O(Y/G)$ of regular functions is a torsion-free $\text{\bf Z}_{2}$-module, since it is a subring of the torsion-free $\text{\bf Z}_{2}$-module $O(Y)$; so $Y/G$ is flat over $\text{\bf Z}_{2}$. The fixed point scheme of $G$ on $Y$ is defined by: $I(x)=(1-x^{2})/x$, and $I(y)=(1-y^{2})/y$, and $I(i)=\sigma(i)-i=-2i$, hence by $\{x^{2}=1,y^{2}=1,2i=0\}$. Together with the equation $i^{2}=-1$ on $Y$, these equations imply set-theoretically that $2=0$, $x=1$, $y=1$, and $i=1$; so the fixed point set of $G$ is a single closed point $P$ in $Y$, with residue field $\text{\bf F}_{2}$. Since $Y$ is regular, it follows that $Y/G$ is regular outside the image of $P$, which we also call $P$.
@@
Fogarty showed that for $G=\text{\bf Z}/p$ acting with an isolated fixed point on a normal scheme $W$ over $\text{\bf F}_{p}$ of dimension at least 3, $W/G$ is not Cohen-Macaulay [12, Proposition 4]. When $W$ has mixed characteristic $(0,p)$, he needed $\dim(W)\geq 4$ to get the same conclusion. Nonetheless, we can build on his ideas to study the 3-dimensional scheme $Y$ in mixed characteristic.
@@
It remains to show that $Y/G$ is terminal. Let us recall the definition. For a normal quasi-projective scheme $X$ over a regular base scheme $S$, Hartshorne defined the canonical sheaf $\omega_{X/S}$ [18, Definition 1.6]. It is a reflexive sheaf of rank 1, or equivalently the sheaf associated to a Weil divisor. In this paper, $S$ will be Spec of the $p$-adic integers or of a field, and we write $K_{X}$ for $\omega_{X/S}$. Toward the end of the proof of Theorem 5.1, we compute $K_{X}$ directly from the definition in our example.
@@
with $a_{j}>0$ for every exceptional divisor $E_{j}$ of $\pi$. If $X$ has a resolution of singularities, terminality of $X$ is equivalent to positivity of the discrepancies $a_{j}$ on this one resolution [18, Corollary 2.12].
Let $Y_{0}=Y$. To prove that our example $Y_{0}/G$ is terminal, one approach would be to construct an explicit resolution of singularities. As with the analogous example in characteristic 2 [26, Theorem 5.1], this can be done by making $G$-equivariant blow-ups of $Y_{0}$ along regular closed subschemes. Namely, we can make $G$-equivariant blow-ups $Y_{2}\to Y_{1}\to Y_{0}$ such that $Y_{2}/G$ is regular. That resolution of $Y_{0}/G$ has dual complex a star, with one edge from a vertex $F_{0}$ to each of 7 other vertices $F_{1},\ldots,F_{7}$.
@@
The fixed point scheme $Y_{1}^{G}$ is defined by: $I(x_{0})=\sigma(x_{0})-x_{0}=x_{0}^{2}(-1-x_{0}y_{2}^{2}+x_{0}^{2}y_{2}^{3})/(1+x_{0})$, $I(y_{1})=x_{0}y_{1}(1-y_{1})$, and $I(y_{2})=x_{0}y_{2}(-1-y_{2}+x_{0}y_{2}+x_{0}y_{2}^{2})$. We know that $Y_{1}^{G}$ (as a set) is contained in $E_{0}$. To focus on the fixed point scheme near $E_{0}$, we can say (more simply): $I(x_{0})=x_{0}^{2}(1+O(x_{0}))$, $I(y_{1})=x_{0}y_{1}(1+y_{1}+O(x_{0}))$, and $I(y_{2})=x_{0}y_{2}(1+y_{2}+O(x_{0}))$. We see that the fixed point scheme $Y_{1}^{G}$ is generically the Cartier divisor $E_{0}=\{x_{1}=0\}$. The bad locus in $E_{0}$ (where that fails) is given by removing a factor of $x_{0}$ from the equations and setting $x_{0}=0$, so we have: $0=x_{0}$, $0=y_{1}(1+y_{1})$, and $0=y_{2}(1+y_{2})$. (Note that $x_{1}=0$ implies $2=0$, by the equation for $U_{0}$.) So the fixed point scheme $Y_{1}^{G}$ (in this chart) is $E_{0}$ as a Cartier divisor except at the 4 points $(x_{0},y_{1},y_{2})$ equal to $(0,0,0)$, $(0,1,0)$, $(0,0,1)$, or $(0,1,1)$.
@@
One further $G$-equivariant blow-up at each of these 7 points suffices to resolve $Y_{1}/G$, but the equations for these blow-ups are a bit messy. Instead, we will use Theorem 2.2 to show that the 7 singular points of $Y_{1}/G$ are all mixed characteristic analogs of the singularity $A^{3}/\mu_{2}$, the simplest terminal singularity whose canonical class is not Cartier. More precisely, each of these 7 singular points is of the form $\frac{1}{2}(1,1,1)$, meaning that it can be written as $Q/\mu_{2}$ for some regular scheme $Q$ of dimension 3 (at an isolated fixed point of $\mu_{2}$). As a result, $Y_{1}/G$ is terminal. With one last calculation, we will deduce that $Y_{0}/G$ is terminal.
@@
Write $K_{Y_{0}}$ for the canonical sheaf $\omega_{Y_{0}/\text{\bf Z}_{2}}$. Since $Y_{0}$ is regular, $K_{Y_{0}}$ is a line bundle, described as follows [18, Definition 1.6]. First, let $R=\text{\bf Z}_{2}[i]/(i^{2}+1)$. Then we have an embedding $D=\operatorname{Spec}R\subset A^{1}_{\text{\bf Z}_{2}}$; write $I$ for the ideal $(i^{2}+1)\subset\text{\bf Z}_{2}[i]$ defining this subscheme. Then the adjunction formula $K_{D}=(K_{A^{1}}+D)|_{D}$ is made into a definition:
@@
In these terms, one trivializing section of $\omega_{R/\text{\bf Z}_{2}}$ is $\alpha:=\frac{di}{i}\cdot f$, where $f\colon I/I^{2}\to R=\text{\bf Z}_{2}[i]/I$ is the map sending $i^{2}+1$ to 1. (Formally, one could think of this section of $\omega_{R/\text{\bf Z}_{2}}$ as $\frac{1}{i}\frac{di}{d(i^{2}+1)}$.) Next, since $\pi\colon Y_{0}\to\operatorname{Spec}R$ is smooth of relative dimension 2, we have $K_{Y_{0}}=\omega_{Y_{0}/\text{\bf Z}_{2}}=\Omega^{2}_{Y_{0}/R}\otimes\pi^{*}\omega_{R/\text{\bf Z}_{2}}$. So one trivializing section of $K_{Y_{0}}$ is $\beta:=\frac{dx}{x}\wedge\frac{dy}{y}\wedge\frac{di}{i}\cdot f$. I claim that this section is fixed by $G$. Indeed, if we extend the action of $G$ on $R$ to $A^{1}_{\text{\bf Z}_{2}}$ by $\sigma(i)=-i$, then $\sigma(f)=f$ and $\sigma(di)=-di$, so $\sigma(\frac{di}{i}f)=\frac{di}{i}f$. Also, $\sigma(dx/x)=-dx/x$ and $\sigma(dy/y)=-dy/y$, from which we see that $\sigma(\beta)=\beta$ as claimed. It follows that the divisor class $K_{Y_{0}/G}$ is linearly equivalent to zero, in particular Cartier. Here $K_{Y_{0}/G}$ is the canonical sheaf in the sense of [18, Definition 1.6]; $Y_{0}/G$ is not Gorenstein, since (as we have shown) it is not Cohen-Macaulay.

### v2

@@ 3 Review of ramification theory
Let $G=\text{\bf Z}/p=\langle\sigma:\sigma^{p}=1\rangle$ act nontrivially on a normal noetherian integral scheme $Y$. Assume that $Y$ is of finite type over a field or over $\text{\bf Z}_{p}$, so that we can talk about the canonical class $K_{Y}$. Write $f$ for the quotient map $Y\rightarrow Y/G$. For each irreducible divisor $E$ in $Y$ that is mapped into itself by $G$, let $F$ be its image (as an irreducible divisor) in $Y/G$. Assume that $p=0$ on $E$. There are several invariants we want to compute in this situation: the ramification index of the divisor $E$ in $Y$ (the positive integer $e$ such that $f^{*}F=eE$), and the coefficient $c$ of $E$ in the ramification divisor (meaning that $K_{Y}=f^{*}K_{Y/G}+cE$ near the generic point of $E$). Another name for $c$ is the valuation of the different $v_{E}(\mathcal{D}_{k(Y)/k(Y/G)})$, where the valuation $v_{E}$ on the function field $k(Y)$ is the order of vanishing along $E$. Here $ef=p$, where $f$ is the degree of the field extension $k(E)$ over $k(F)$.
@@
Define the Artin ramification number $i(E)$ of $Y$ over $Y/G$ along $E$ to be the coefficient of $E$ in the fixed point scheme $Y^{G}$. Equivalently, in terms of $I(a)=\sigma(a)-a$:
@@
In the ramified case, the field $k(E)$ is purely inseparable over $k(F)$, with degree $f$ equal to 1 or $p$. We can distinguish the two cases as follows. Since $ef=p$, either $e=p$ and $f=1$ (called wild ramification) or $e=1$ and $f=p$ (called fierce ramification). Let $t$ be a defining function of $E$, that is, a rational function on $Y$ with valuation $v_{E}(t)=1$. It is clear that $v_{E}(I(t))\geq i(E)$. A very convenient criterion is: $Y\to Y/G$ is wildly ramified along $E$ if and only if equality holds [30, section 2.1]. Otherwise, it is fiercely ramified.
Furthermore, we can compute the relative canonical class (that is, the valuation of the different) as follows, correcting a typo in [30, section 2.1].
Lemma 3.1.
The valuation of the different is $(p-1)i(E)$. When $Y$ is regular, so that $Y/G$ is Q-factorial, we can equivalently say that
@@
Proof.
If $G$ acts nontrivially on $E$ (the unramified case), then $i(E)=0$ and the statement is clear. So assume that $G$ fixes $E$. The local ring $O_{Y,E}$ is a discrete valuation ring. The algebra $O_{Y,E}$ is generated by one element $y$ as an algebra over $O_{Y/G,F}$; one can take $y$ to be a uniformizer in $O_{Y,E}$ if $E$ is wildly ramified, and an element of $O_{Y,E}$ whose restriction to $k(E)$ is not in $k(F)$ if $E$ is fiercely ramified [30, section 2.1]. By Lemma 2.3, we have $i(E)=v_{E}(I(y))$.
Let $u(X)$ be the minimal polynomial of $y$ over $k(Y/G)$. In this situation of a monogenic algebra extension, the different $\mathcal{D}_{k(Y)/k(Y/G)}$ is generated by $u^{\prime}(y)$ [26, III, Corollary 2 to Proposition 11]. But $u(X)=\prod_{j=0}^{p-1}(X-\sigma^{j}(y))$. So
$u^{\prime}(y)=\prod_{j=1}^{p-1}(y-\sigma^{j}(y)),$
and hence
$v_{E}(\mathcal{D}_{k(Y)/k(Y/G)})=v_{E}(u^{\prime}(y))=(p-1)i(E),$
using that $i(E)$ is unchanged if we replace $\sigma$ by another generator of $G$. ∎
@@ 4 Toric divisors
    Let $G=\text{\bf Z}/p$ act on a regular scheme $U$, with the assumptions of Theorem 2.2. Assume moreover that the function $e$ is the greatest common divisor of the functions $I(x_{i})$ for $i=1,\ldots,n$ in the local ring $O_{U,P}$. Then, for each $y\in\operatorname{\mathfrak{m}}_{U}$ such that $\{y=0\}$ is an irreducible divisor $E\subset U$ and $I(y)\in(ey)$, $E$ is the pullback of a Weil divisor $F$ in $U/G$.
@@
    Continue to assume that $I(x_{i})\in e\operatorname{\mathfrak{m}}_{U}$ for $i=1,\ldots,n$, so that $U/G$ is the quotient of a regular scheme $Q$ by $\mu_{p}$. Let $y_{1},\ldots,y_{r}$ be functions on $U$, vanishing at $P$, that are linearly independent in $\operatorname{\mathfrak{m}}_{U}/\operatorname{\mathfrak{m}}_{U}^{2}$. Suppose that $I(y_{j})\in(ey_{j})$ for $j=1,\ldots,r$. Then the corresponding $\mu_{p}$-eigenfunctions $h_{1},\ldots,h_{r}$ on $Q$ (from (2)) are linearly independent in $\operatorname{\mathfrak{m}}_{Q}/\operatorname{\mathfrak{m}}_{Q}^{2}$. That is, these functions are part of a toric coordinate system on $Q$. Finally, if $I(s)=es$ in Theorem 2.2 (as we can assume), then the $\mu_{p}$-weight of $h_{j}$ is equal to the eigenvalue of $\varphi(y):=I(y)/e$ on $y_{j}\in\operatorname{\mathfrak{m}}_{U}/\operatorname{\mathfrak{m}}_{U}^{2}$, which is in $\text{\bf F}_{p}$.
@@
(1) Let $E$ be the irreducible divisor $\{y=0\}$ in $U$. The assumptions imply that $E$ is mapped into itself by $G$, and that there is an $i\in\{1,\ldots,n\}$ such that $v_{E}(I(y))>v_{E}(I(x_{i}))$. By section 3, $f\colon U\to U/G$ is either unramified (if $E$ is not fixed by $G$) or fiercely ramified along $E$. In either case, there is an irreducible divisor $F$ on $U/G$ such that $E=f^{*}F$, as we want. (The divisor $F$ need not be Cartier, but the pullback of a Weil divisor is still a Weil divisor (with integer coefficients). Indeed, $F$ is a Cartier divisor outside a codimension-2 subset of $U/G$, by normality of $U/G$.)
@@
(2) Since $Q$ is regular, the pullback of $F$ to $Q$ is a Cartier divisor, hence (after shrinking $U$ and $Q$ around $P$) of the form $\{t=0\}$ for some function $t\in\operatorname{\mathfrak{m}}_{Q}-\{0\}$. Clearly the divisor $\{t=0\}$ is $\mu_{p}$-invariant. I claim that $t$ times some unit is a $\mu_{p}$-eigenfunction on $Q$. Indeed, in algebraic terms, the action of $\mu_{p}$ on $Q$ makes $O(Q)$ a comodule over $O(\mu_{p})$, and the ideal $(t)$ is an sub-$O(\mu_{p})$-comodule. Every $O(\mu_{p})$-comodule (with no finiteness assumption needed) is the direct sum of its weight spaces, indexed by $\text{\bf Z}/p$. So we can write $t=t_{0}+\cdots+t_{p-1}$ with $t_{i}\in(t)$ and $t_{i}$ of weight $i$. Since $t_{i}\in(t)$, we can write $t_{i}=a_{i}t$ for some $a_{i}\in O_{Q,P}$. Since $O_{Q,P}$ is regular, it is a domain, and hence $1=a_{0}+\cdots+a_{p-1}$. So at least one $a_{i}$ is not in $\operatorname{\mathfrak{m}}_{Q}$, hence is a unit. Then $h:=t_{i}=a_{i}t$ is a unit times $t$ and also a $\mu_{p}$-eigenfunction (of weight $i$), as we want.
@@
(3) After multiplying $e$ by a unit, we can assume that $I(s)=es$, in the terminology of Theorem 2.2. The assumption that $I(y_{j})\in(ey_{j})$ for $j=1,\ldots,r$ implies that $y_{1},\ldots,y_{r}$ in $\operatorname{\mathfrak{m}}_{U}/\operatorname{\mathfrak{m}}_{U}^{2}$ are eigenvectors of the map $\varphi$. By the proof of Theorem 2.2, the corresponding eigenvalues are in $\text{\bf F}_{p}\subset k_{U}$.
@@
For $j=1,\ldots,r$, we know from (1) that the divisor $\{y_{j}=0\}$ on $U$ is pulled back from a divisor $F_{j}$ on $U/G$. (Here $F_{j}$ is a Weil divisor, but it is a Cartier divisor outside a codimension-2 subset of $U/G$, since $U/G$ is normal.) By (2), $F_{j}$ pulls back to a divisor $\{h_{j}=0\}$ on $Q$ with $h_{j}$ a $\mu_{p}$-eigenfunction. It follows that the Cartier divisors $\{y_{j}=0\}$ on $U$ and $\{h_{j}=0\}$ on $Q$ have the same pullback to $W$; that is, $h_{j}=y_{j}(\text{unit})$ on $W$. Since $h_{j}$ is a $\mu_{p}$-eigenfunction of some weight $b_{j}\in\{0,\ldots,p-1\}$, we have $h_{j}=g_{j}w^{b_{j}}$ for some function $g_{j}$ on $U$, in the notation of the proof of Theorem 2.2. Therefore, $g_{j}=y_{j}(\text{unit})$ on $U$. Since $y_{1},\ldots,y_{r}$ are linearly independent in $\operatorname{\mathfrak{m}}_{U}/\operatorname{\mathfrak{m}}_{U}^{2}$, the same is true for $g_{1},\ldots,g_{r}$. By the proof of Theorem 2.2, the $\mu_{p}$-weight of $h_{j}$ is equal to the eigenvalue of $\varphi$ on the eigenvector $g_{j}\in\operatorname{\mathfrak{m}}_{U}/\operatorname{\mathfrak{m}}_{U}^{2}$. Since $g_{j}\in\operatorname{\mathfrak{m}}_{U}/\operatorname{\mathfrak{m}}_{U}^{2}$ is a nonzero multiple of $y_{j}$, this is the same as the eigenvalue of $\varphi$ on $y_{j}$.
@@ 5 The example over the 2-adic integers
The scheme $Y$ is regular, being an open subset of the affine plane over the discrete valuation ring $\text{\bf Z}_{2}[\zeta_{4}]=\text{\bf Z}_{2}[i]/(i^{2}+1)$. Since $Y$ is a normal integral affine scheme of dimension 3, so is $Y/G$. The ring $O(Y/G)$ of regular functions is a torsion-free $\text{\bf Z}_{2}$-module, since it is a subring of the torsion-free $\text{\bf Z}_{2}$-module $O(Y)$; so $Y/G$ is flat over $\text{\bf Z}_{2}$. The fixed point scheme of $G$ on $Y$ is defined by: $I(x)=(1-x^{2})/x$, $I(y)=(1-y^{2})/y$, and $I(i)=\sigma(i)-i=-2i$, hence by $\{x^{2}=1,y^{2}=1,2i=0\}$. Together with the equation $i^{2}=-1$ on $Y$, these equations imply set-theoretically that $2=0$, $x=1$, $y=1$, and $i=1$; so the fixed point set of $G$ is a single closed point $P$ in $Y$, with residue field $\text{\bf F}_{2}$. Since $Y$ is regular, it follows that $Y/G$ is regular outside the image of $P$, which we also call $P$.
@@
Fogarty showed that for $G=\text{\bf Z}/p$ acting with an isolated fixed point on a normal scheme $W$ over $\text{\bf F}_{p}$ of dimension at least 3, $W/G$ is not Cohen-Macaulay [13, Proposition 4]. When $W$ has mixed characteristic $(0,p)$, he needed $\dim(W)\geq 4$ to get the same conclusion. Nonetheless, we can build on his ideas to study the 3-dimensional scheme $Y$ in mixed characteristic.
@@
It remains to show that $Y/G$ is terminal. Let us recall the definition. For a normal quasi-projective scheme $X$ over a regular base scheme $S$, Hartshorne defined the canonical sheaf $\omega_{X/S}$ [20, Definition 1.6]. It is a reflexive sheaf of rank 1, or equivalently the sheaf associated to a Weil divisor. In this paper, $S$ will be Spec of the $p$-adic integers or of a field, and we write $K_{X}$ for $\omega_{X/S}$. Toward the end of the proof of Theorem 5.1, we compute $K_{X}$ directly from the definition in our example.
@@
with $a_{j}>0$ for every exceptional divisor $E_{j}$ of $\pi$. If $X$ has a resolution of singularities, terminality of $X$ is equivalent to positivity of the discrepancies $a_{j}$ on this one resolution [20, Corollary 2.12].
Let $Y_{0}=Y$. To prove that our example $Y_{0}/G$ is terminal, one approach would be to construct an explicit resolution of singularities. As with the analogous example in characteristic 2 [29, Theorem 5.1], this can be done by making $G$-equivariant blow-ups of $Y_{0}$ along regular closed subschemes. Namely, we can make $G$-equivariant blow-ups $Y_{2}\to Y_{1}\to Y_{0}$ such that $Y_{2}/G$ is regular. That resolution of $Y_{0}/G$ has dual complex a star, with one edge from a vertex $F_{0}$ to each of seven other vertices $F_{1},\ldots,F_{7}$.
@@
The fixed point scheme $Y_{1}^{G}$ is defined by: $I(x_{0})=\sigma(x_{0})-x_{0}=x_{0}^{2}(-1-x_{0}y_{2}^{2}+x_{0}^{2}y_{2}^{3})/(1+x_{0})$, $I(y_{1})=x_{0}y_{1}(1-y_{1})$, and $I(y_{2})=x_{0}y_{2}(-1-y_{2}+x_{0}y_{2}+x_{0}y_{2}^{2})$. We know that $Y_{1}^{G}$ (as a set) is contained in $E_{0}$. To focus on the fixed point scheme near $E_{0}$, we can say (more simply): $I(x_{0})=x_{0}^{2}(1+O(x_{0}))$, $I(y_{1})=x_{0}y_{1}(1+y_{1}+O(x_{0}))$, and $I(y_{2})=x_{0}y_{2}(1+y_{2}+O(x_{0}))$. We see that the fixed point scheme $Y_{1}^{G}$ is generically the Cartier divisor $E_{0}=\{x_{0}=0\}$. The bad locus in $E_{0}$ (where that fails) is given by removing a factor of $x_{0}$ from the equations and setting $x_{0}=0$, so we have: $0=x_{0}$, $0=y_{1}(1+y_{1})$, and $0=y_{2}(1+y_{2})$. (Note that $x_{0}=0$ implies $2=0$, by the equation for $U_{0}$.) So the fixed point scheme $Y_{1}^{G}$ (in this chart) is $E_{0}$ as a Cartier divisor except at the 4 points $(x_{0},y_{1},y_{2})$ equal to $(0,0,0)$, $(0,1,0)$, $(0,0,1)$, or $(0,1,1)$.
@@
One further $G$-equivariant blow-up at each of these 7 points suffices to resolve $Y_{1}/G$, but the equations for these blow-ups are a bit messy. Instead, we will use Theorem 2.2 to show that the 7 singular points of $Y_{1}/G$ are all mixed-characteristic analogs of the singularity $A^{3}/\mu_{2}$, the simplest terminal singularity whose canonical class is not Cartier. More precisely, each of these 7 singular points is of the form $\frac{1}{2}(1,1,1)$, meaning that it can be written as $Q/\mu_{2}$ for some regular scheme $Q$ of dimension 3 (at an isolated fixed point of $\mu_{2}$). As a result, $Y_{1}/G$ is terminal. With one last calculation, we will deduce that $Y_{0}/G$ is terminal.
@@
Write $K_{Y_{0}}$ for the canonical sheaf $\omega_{Y_{0}/\text{\bf Z}_{2}}$. Since $Y_{0}$ is regular, $K_{Y_{0}}$ is a line bundle, described as follows [20, Definition 1.6]. First, let $R=\text{\bf Z}_{2}[i]/(i^{2}+1)$. Then we have an embedding $D=\operatorname{Spec}R\subset A^{1}_{\text{\bf Z}_{2}}$; write $I$ for the ideal $(i^{2}+1)\subset\text{\bf Z}_{2}[i]$ defining this subscheme. Then the adjunction formula $K_{D}=(K_{A^{1}}+D)|_{D}$ is made into a definition:
@@
In these terms, one trivializing section of $\omega_{R/\text{\bf Z}_{2}}$ is $\alpha:=\frac{di}{i}\cdot f$, where $f\colon I/I^{2}\to R=\text{\bf Z}_{2}[i]/I$ is the map sending $i^{2}+1$ to 1. (Formally, one could think of this section of $\omega_{R/\text{\bf Z}_{2}}$ as $\frac{1}{i}\frac{di}{d(i^{2}+1)}$.) Next, since $\pi\colon Y_{0}\to\operatorname{Spec}R$ is smooth of relative dimension 2, we have $K_{Y_{0}}=\omega_{Y_{0}/\text{\bf Z}_{2}}=\Omega^{2}_{Y_{0}/R}\otimes\pi^{*}\omega_{R/\text{\bf Z}_{2}}$. So one trivializing section of $K_{Y_{0}}$ is $\beta:=\frac{dx}{x}\wedge\frac{dy}{y}\wedge\frac{di}{i}\cdot f$. I claim that this section is fixed by $G$. Indeed, if we extend the action of $G$ on $R$ to $A^{1}_{\text{\bf Z}_{2}}$ by $\sigma(i)=-i$, then $\sigma(f)=f$ and $\sigma(di)=-di$, so $\sigma(\frac{di}{i}f)=\frac{di}{i}f$. Also, $\sigma(dx/x)=-dx/x$ and $\sigma(dy/y)=-dy/y$, from which we see that $\sigma(\beta)=\beta$ as claimed. It follows that the divisor class $K_{Y_{0}/G}$ is linearly equivalent to zero, in particular Cartier. Here $K_{Y_{0}/G}$ is the canonical sheaf in the sense of [20, Definition 1.6]; $Y_{0}/G$ is not Gorenstein, since (as we have shown) it is not Cohen-Macaulay.